# Artifact-free Sound Quality in DNN-based Closed-loop Systems for Audio Processing

### v2

@@ VI-B Biophysical properties of auditory modules
We employed auditory neuroscience techniques to quantify key properties of the dCoNNear-based auditory elements, comparing them against the reference analytical model and prior CoNNear outputs.
Figures 8 and 9 compare the basilar membrane (BM) excitation patterns and QERB values of the proposed dCoNNear model with the reference TL model and prior CoNNearcochlear. As shown in Fig. 8, dCoNNear exhibited performance comparable to both the prior CoNNear model and the TL model, suggesting that dCoNNear accurately captures the shape and compression properties of pure-tone cochlear excitation patterns. Fig. 9 illustrates a close match between dCoNNear, the reference TL model, and CoNNearcochlear, indicating that dCoNNear reliably replicates the frequency-dependent characteristics of cochlear responses.
@@
Figure 9: Comparison of cochlear filter tuning (QERB) across various center frequencies (CFs) using 100 $\mu$s clicks at 40 and 70 dB SPL among the dCoNNear model, the prior CoNNear model, and the target model.
@@
In conclusion, the dCoNNear-based auditory models faithfully replicate the auditory properties of the reference analytical models, ensuring accurate and reliable performance in auditory signal processing tasks.
@@ VI-C Evaluation of restoration performance
Table LABEL:tab:NRMSE presents the average normalized root-mean-square errors (NRMSEs) between the normal NH and HI AN population responses. These values were computed for the TIMIT utterances before and after processing with the HA models trained using the prior CoNNear and dCoNNear-based systems. To ensure a fair comparison, the same loss function was applied to train both HA models. Both models effectively reduced the NRMSEs of the AN population responses across different input levels, with more effective processing observed for speech presented at lower levels. However, the prior CoNNear exhibited marginally better performance than the dCoNNear. This may be attributed to the tonal artifacts generated by the HA models trained from the prior CoNNear-based system. As noted in [36], temporal peaks in sound can lead to increased AN responses, and high-frequency tonal artifacts can contribute to these elevated AN population responses.
@@
TABLE IV: The average NRMSE percentages across different HA models for input levels between 40 and 70 dB SPL.
40 50 60 70
Unprocessed 29.08 20.13 15.86 14.25
@@
Figure 13: Simulated auditory nerve (AN) population responses for hearing aid (HA) models trained using different frameworks at 70 dB SPL. (a) A clean speech segment is shown before and after processing with models trained using prior CoNNear (blue) and dCoNNear (green) systems. (b) The hearing-impaired (HI) AN population responses were enhanced after processing the input stimulus with different HA models. Ideal compensation would align the HI AN population responses with the NH AN population responses.
VII Discussion
@@
Current closed-loop frameworks suffer from non-ideal sampling processes, which can introduce undesired artifacts and thus degrade the resulting audio quality. In this study, we systematically characterized the artifacts generated by different upsampling methods and examined those associated with transposed convolution within the context of closed-loop, autoencoder-based hearing aid algorithms. These artifacts originated from the interaction between problematic upsampling operations and spectral replicas by the upsampling layers. The artifacts persisted in the closed-loop system after training, leading to tonal artifacts in the resulting audio.
@@
We propose an artifact-free architecture, dCoNNear, designed to integrate into closed-loop systems, exemplified here for hearing aid algorithms. In contrast to the existing autoencoder-based closed-loop system [6], the dCoNNear-based architecture can significantly reduce the artifacts associated with prior CoNNear models, thereby improving the sound quality of the resulting audio. This advancement provides an artifact-free framework for audio applications, delivering high sound quality for future audio processing algorithms.
[…]
The current dCoNNear models are limited to wave-to-wave applications due to the absence of upsampling operations. This restricts their use to specific tasks such as speech enhancement, speech conversion, and hearing aids. Expanding the architecture to include more versatile processing capabilities could broaden its applicability and effectiveness across various audio processing domains.
@@
The use of the Speech-to-Reverberation Modulation Energy Ratio (SRMR) as an evaluation metric provided valuable insights into the sound quality of speech samples. However, SRMR falls short in evaluating other sounds types, such as music and environmental noises. This study suggests the need for broader objective metrics to assess a broader range of audio types. Additionally, while objective metrics are useful, subjective evaluations involving end-users are equally crucial. Incorporating user-centric evaluation metrics, such as the Hearing in Noise Test (HINT) [37] and Multiple Stimuli with Hidden Reference and Anchor (MUSHRA) [38], would provide a more comprehensive understanding of the perceptual benefits and real-world performance of DNN-based audio processing systems.
VIII Conclusion
@@
In this study, we systematically characterized the artifacts generated by different upsampling operations and examined those associated with transposed convolution within the context of closed-loop, autoencoder-based hearing aid algorithms. These artifacts stemmed from the interaction between problematic upsampling operations and spectral replicas created by the upsampling layers, compromising the sound quality of the resulting audio applications. To address these issues, we proposed an artifact-free architecture, dCoNNear, designed to integrate into closed-loop systems, specifically for hearing aid algorithms. The dCoNNear architecture incorporates a series of stacked FIR-like memory blocks, eliminating the need for sampling operations and demonstrating strong capability in modeling long-term dependencies. Our results demonstrated that the dCoNNear-based models could significantly reduce the artifacts associated with autoencoder-based prior CoNNear models and improve the resulting audio quality. Additionally, dCoNNear not only accurately emulates all non-DNN-based biophysical auditory models but also shows comparable compensation performance to prior CoNNear models. The dCoNNear-based closed-loop framework holds great promise for advancing audio technologies, including hearing aids, speech enhancement, and speech conversion, with high sound quality.

### v3

@@ VI-B Biophysical properties of auditory modules
To ensure that changing the CoNNear architecture did not compromise the performance, we employed auditory neuroscience techniques to quantify key properties of the dCoNNear-based auditory elements, comparing them against the reference analytical model and prior CoNNear outputs.
Figures 9 and 10 compare the basilar membrane (BM) excitation patterns and QERB values of the proposed dCoNNear model with the reference TL model and prior CoNNearcochlear. As shown in Fig. 9, dCoNNear exhibited performance comparable to both the prior CoNNear model and the TL model, suggesting that dCoNNear accurately captures the shape and compression properties of pure-tone cochlear excitation patterns. Fig. 10 illustrates a close match between dCoNNear, the reference TL model, and CoNNearcochlear, indicating that dCoNNear reliably replicates the frequency-dependent characteristics of cochlear responses.
@@
Figs. 11 and 12 present the IHC excitation patterns and the half-wave rectified IHC receptor potential, compared to the reference IHC model and prior CoNNearIHC predictions. The excitation patterns from the dCoNNearIHC model closely align with those of both the reference IHC model and prior CoNNearIHC (Fig. 11). Fig. 12 shows the RMS of the half-wave rectified IHC receptor potential, $V\textsubscript{IHC}$, in response to a 4-kHz pure tone across input levels ranging from 0 to 100 dB SPL. The blue curve, representing the dCoNNearIHC model, follows the trends of the black and orange curves corresponding to the reference and prior models, respectively. This consistency indicates a nearly linear relationship with SPL up to 90 dB, followed by compressive growth beyond 90 dB SPL. These results suggest that the dCoNNearIHC model accurately captures the potential-level growth characteristics.
@@
In conclusion, the dCoNNear-based auditory models faithfully replicate the auditory properties of the reference analytical models, ensuring accurate and reliable performance in auditory signal processing tasks even with the new dCoNNear architecture.
@@ VI-C Evaluation of restoration performance
Table LABEL:tab:NRMSE presents the average normalized root-mean-square errors (NRMSEs) between the normal NH and HI AN population responses for utterances from the TIMIT and LibriTTS corpora. These values were calculated for unprocessed and processed signals using hearing-aid (HA) models trained with either the prior CoNNear or the proposed dCoNNear frameworks. To ensure a fair comparison, both HA models were trained using the same loss function and optimization settings. In both corpora, the results followed consistent trends. Both the prior CoNNear and dCoNNear-based HA models effectively reduced the NRMSE across input levels ranging from 40 to 70 dB SPL, with the largest improvements observed at lower input levels (40–50 dB SPL), where HA processing plays a more significant role. However, the prior CoNNear exhibited marginally better performance than the dCoNNear. This may be attributed to the tonal artifacts generated by the HA models trained from the prior CoNNear-based system. As noted in [drakopoulos2022model], temporal peaks or spurious high-frequency tones can result in increased AN responses, potentially narrowing the gap between impaired and reference NH responses in a way that lowers NRMSE but does not reflect true restoration.
@@
VI-D Real-Time Inference Evaluation
To assess the real-time performance of the proposed audio processor, we measured the inference latency on an Intel i7-1265U CPU and an NVIDIA A30 GPU. A total of 100 speech samples from the test set were processed, each consisting of 512 samples (equivalent to 25.6 ms at a 20 kHz sampling rate). The dCoNNear-based HA model achieved average processing times of 16.42 ms on the CPU and 3.81 ms on the GPU, corresponding to real-time factors (RTFs) of 0.64 and 0.15, respectively, indicating real-time capability on both platforms. For comparison, the previous CoNNear-based HA model from [drakopoulos2023neural] processed the same input frames in 3.98 ms on the CPU and 1.57 ms on the GPU, yielding RTFs of 0.15 and 0.06, respectively. While the dCoNNear-based model introduces additional computational complexity compared to the original CoNNear model, it still meets real-time constraints, even on a standard CPU.
TABLE V: Average NRMSE (%) across different HA models for input levels between 40 and 70 dB SPL, evaluated on the TIMIT and LibriTTS corpora.
Dataset Model 40 50 60 70
TIMIT Unprocessed 29.08 20.13 15.86 14.25
@@ VI-D Real-Time Inference Evaluation
Figure 14: Simulated auditory nerve (AN) population responses for hearing-aid (HA) models trained using different frameworks at 70 dB SPL. (a) A clean speech segment is shown before and after processing with models trained using prior CoNNear (blue) and dCoNNear (green) systems. (b) The hearing-impaired (HI) AN population responses were enhanced after processing the input stimulus with different HA models. Ideal compensation would align the HI AN population responses with the NH AN population responses.
Figure 15: The framework of training speech enhancement models. The auditory modules are based on prior CoNNear and dCoNNear.
VII Speech Enhancement Application
Beyond its primary validation in hearing-aid design, the dCoNNear framework’s inherent ability to generate artifact-free audio extends its utility to various wave-to-wave audio processing tasks. This section, therefore, presents a proof-of-concept experiment to demonstrate dCoNNear’s efficacy in a closed-loop speech enhancement (SE) task, illustrating its broader applicability.
VII-A Framework
The training framework for the speech enhancement (SE) model is illustrated in Fig. 15. Both processing pathways include auditory modules that simulate normal-hearing auditory profiles. In the reference pathway, clean speech is fed directly into the auditory system. In the parallel pathway, the speech enhancement model first processes a noisy input, and its output is then passed through the same auditory modules. Both the auditory modules and the SE model follow the same architectural design as those used in the hearing-aid application described earlier.
VII-B Training
We used utterances from the INTERSPEECH 2021 DNS Challenge dataset [reddy2021interspeech] for training and utterances from LibriTTS [zen2019libritts] for testing. For noise data, Audioset [gemmeke2017audio] was used during training, while Freesound [fonseca2017freesound] served as the test noise source. Noisy speech samples were generated by randomly mixing clean utterances with noise segments at signal-to-noise ratios (SNRs) ranging from –5 dB to 5 dB in 1 dB increments. This resulted in 100 hours of training data and 5 hours of validation data. A separate 1-hour test set was created at each of the three SNR levels: –5 dB, 0 dB, and 5 dB. All samples were resampled to 20 kHz to match the frequency range of the auditory models and normalized to 70 dB SPL to ensure consistency with the closed-loop system.
The SE model was trained to minimize the mean absolute error (MAE) between the outputs of the two auditory pathways. During training, the parameters of the auditory modules were kept frozen. For comparison, we also trained a prior CoNNear-based SE model using the same training procedure and dataset. This allowed for a direct evaluation of the benefits introduced by the dCoNNear architecture in a closed-loop speech enhancement task.
VII-C Results
We evaluated the sound quality of the prior CoNNear- and dCoNNear-based speech enhancement (SE) models using three objective metrics: DNSMOS (overall) [reddy2021dnsmos], PESQ [rix2001perceptual], and ESTOI [jensen2016algorithm]. Table VI presents the performance comparison across input SNRs of –5 dB, 0 dB, and 5 dB. Across all SNR conditions, the dCoNNear-based SE model consistently achieved higher scores in DNSMOS, PESQ, and ESTOI, indicating improved perceptual quality and intelligibility over the prior CoNNear-based model.
To further analyze the presence of artifacts, Fig. 16 displays spectrograms of a representative speech example corrupted by babble noise and subsequently enhanced by the two SE models. The output of the prior CoNNear-based model exhibited distinct tonal artifacts, visible as horizontal lines in the spectrogram, which were not present in the original noisy input. In contrast, the spectrogram from the dCoNNear-based model showed a marked reduction in these artifacts. This qualitative evidence supported the objective results and demonstrated that dCoNNear effectively reduces structural artifacts in the closed-loop speech enhancement task. The resulting audio samples are provided online  [^2] https://github.com/chuan997/Trans24_audio_samples.
The results confirm the effectiveness of dCoNNear as a robust and generalizable solution for artifact-free closed-loop audio processing, extending its applicability beyond hearing-aid processing to speech enhancement tasks and potentially other wave-to-wave audio domains.
TABLE VI: Comparison of different models in terms of various objective metrics. Higher values indicate better performance and bold fonts highlight the best performance.
Metrics Models SNRs (dB)
-5 0 5
DNSMOS Unprocessed 1.08 1.29 1.5
prior CoNNear 1.57 2.32 2.54
dCoNNear 2.87 3.01 3.22
PESQ Unprocessed 1.04 1.27 1.45
prior CoNNear 1.18 1.44 1.65
dCoNNear 1.3 1.63 2.12
ESTOI Unprocessed 0.35 0.48 0.6
prior CoNNear 0.39 0.59 0.66
dCoNNear 0.6 0.74 0.82
Figure 16: Spectrograms of (a) noisy speech at 0 dB SNR with babble noise, (b) enhancement by prior CoNNear-based model, (c) enhancement by dCoNNear-based model, and (d) clean speech.
VIII Discussion
@@
Current closed-loop frameworks suffer from non-ideal sampling processes, which can introduce undesired artifacts when placed in closed-loop frameworks and can degrade the audio quality of the resulting application. In this study, we systematically characterized the artifacts generated by different upsampling methods and examined those associated with transposed convolution within the context of closed-loop, autoencoder-based hearing-aid algorithms. These artifacts originated from the interaction between problematic upsampling operations and spectral replicas by the upsampling layers. The artifacts persisted in the closed-loop system after training, leading to tonal artifacts in the resulting audio.
@@
We propose an artifact-free architecture, dCoNNear, designed to integrate into closed-loop systems, exemplified here for hearing-aid algorithms. In contrast to the existing autoencoder-based closed-loop system [drakopoulos2023neural], the dCoNNear-based architecture can significantly reduce the artifacts associated with prior CoNNear models, thereby improving the sound quality of the resulting audio. This advancement provides an artifact-free framework for audio applications, delivering high sound quality for future audio processing algorithms.
From the perspective of the human auditory system, dCoNNear’s modifications significantly enhance fidelity by eliminating DNN architectural artifacts, such as tonal and aliasing distortions, which are inherently absent in natural human auditory processing.These artifacts can compromise the accuracy of bio-inspired models and distort perceptual outcomes. By removing them, dCoNNear enables a more faithful simulation of auditory function, resulting in artifact-free representations. This is particularly important for applications such as personalized hearing-aid algorithms, where the goal is to restore simulated auditory-nerve responses in hearing-impaired models to closely match those of normal-hearing references. Experimental results confirm that dCoNNear accurately models advanced biophysical stages without introducing spurious artifacts, achieving stronger alignment with the functional properties of the human auditory pathway.
While dCoNNear-based models successfully minimize artifacts associated with prior CoNNear, the SRMR scores of the processed samples decreased relative to the inputs, as shown in Table III and IV. One potential reason is the compressive amplification characteristics of the trained HA model, where quieter sounds are amplified more than louder sounds. This process amplifies low-level noise, which subsequently degrades sound quality. Therefore, incorporating a noise reduction module should be considered in future works. Another contributing factor could be the training process, which aimed to minimize differences between NH and HI auditory processing but may introduce non-linear distortions that affect sound quality. This study underscores the need for future research focused on developing more optimal loss functions or post-processing filters. Such developments would enhance the perceptual benefits of audio processing systems.
@@
The use of the Speech-to-Reverberation Modulation Energy Ratio (SRMR) as an evaluation metric provided valuable insights into the sound quality of speech samples. However, SRMR falls short in evaluating other sounds types, such as music and environmental noises. This study suggests the need for broader objective metrics to assess a broader range of audio types.
While user-centric evaluation methods such as the Hearing in Noise Test (HINT) [37] and Multiple Stimuli with Hidden Reference and Anchor (MUSHRA) [38] offer valuable insights into perceptual experience, the present work focuses on addressing architectural artifacts in DNN-based audio models. We therefore adopt established objective metrics that correlate with perceived quality and intelligibility. Given the proof-of-concept nature of this work, integrating formal listening tests remains a valuable direction for future research, particularly as the system advances toward real-world deployment.
IX Conclusion
@@
In this work, we proposed dCoNNear, a novel architecture specifically designed to suppress spurious artifacts—most notably tonal and aliasing artifacts introduced by non-ideal downsampling and upsampling layers. We demonstrated its effectiveness in closed-loop applications for both hearing-aid algorithms and speech enhancement. The dCoNNear architecture incorporates a series of stacked FIR-like memory blocks, eliminating the need for sampling operations and demonstrating strong capability in modeling long-term dependencies. Our results showed that dCoNNear-based models significantly reduced the artifacts associated with prior autoencoder-based CoNNear systems and improved perceptual sound quality across both tasks. Furthermore, dCoNNear not only accurately emulated all non-DNN-based biophysical auditory models but also achieved comparable compensation performance to the prior CoNNear in hearing-aid processing and outperformed it in speech enhancement. Overall, the dCoNNear-based closed-loop framework holds great promise for advancing audio technologies, including hearing aids and speech enhancement with high sound quality.